OpenStudy (hannahh<3):

Hey guys!! Plz help me!!! t/f: When solving inequalities, the addition and subtraction of integers to not affect the inequality sign. A friend of mine says true, but I'm pretty posotive that it's false. Help?

OpenStudy (austinl):

Addition and subtraction doesn't change the sign at all, However, multiplication or division of a negative will flip the sign, \(<\rightarrow>, >\rightarrow<~\&~\le\rightarrow\ge,\ge\rightarrow\le\) Make sense?

OpenStudy (hannahh<3):

but if you do 4-6, you get -2, changing the integer sighn, or -5+6=1, also changing the integer sign.

OpenStudy (austinl):

Lets say for instance you have, \(6x-4>8\) What would \(x\) be?

OpenStudy (hannahh<3):

x>2

OpenStudy (austinl):

Correct, now lets change it slightly. \(-6x-4>8\)

OpenStudy (hannahh<3):

it would still be two because 12 is bigger than 6.

OpenStudy (austinl):

Incorrect, mostly. \(-6x-4>8\) \(-6x>12\) Now, we are dividing by a negative six(-6). That means that we flip the sign. \(\displaystyle \frac{-6x}{-6}<\frac{12}{-6}\) \(x<-2\)

OpenStudy (hannahh<3):

oh. so the answer would be true. I get it, thanks austin!!!

OpenStudy (austinl):

Indeed :) You are very welcome!